# BEVRender: Vision-based Cross-view Vehicle Registration in Off-road GNSS-denied Environment

### v1

@@ Abstract
We introduce BEVRender, a novel learning-based approach for the localization of ground vehicles in Global Navigation Satellite System (GNSS)-denied off-road scenarios. These environments are typically challenging for conventional vision-based state estimation due to the lack of distinct visual landmarks and the instability of vehicle poses. To address this, BEVRender generates high-quality local bird’s eye view (BEV) images of the local terrain. Subsequently, these images are aligned with a geo-referenced aerial map via template-matching to achieve accurate cross-view registration. Our approach overcomes the inherent limitations of visual inertial odometry systems and the substantial storage requirements of image-retrieval localization strategies, which are susceptible to drift and scalability issues, respectively. Extensive experimentation validates BEVRender’s advancement over existing GNSS-denied visual localization methods, demonstrating notable enhancements in both localization accuracy and update frequency. The code for BEVRender will be made available soon.
@@ III Methodology
Our system contains three main components: a feature encoder mapping visual representation from camera to top-down view, a rendering head to decode learned features and render top-down BEV images, and a image registration component for localization. An overview of our system is shown in Fig. 1.
@@
$\mathcal{X}_{t}=\Big{[}x_{t},y_{t},\theta_{t}\Big{]}\in\mathbf{SE}(2).$ (1)
We assume the prediction for azimuth angle $\theta_{t}$ from VIO is accurate but the position estimates ($x_{t}$ and $y_{t}$) may drift over time. Our system aggregates consecutive frames to construct a top-down representation of the environment for map registration.
We define a 3D BEV space centered on the vehicle with a length of $L$, a width of $W$, and a height of $H$. The space is divided into $l\times w\times h$ grid cells, so that each cell represents a cubic size of $\frac{L}{l}\times\frac{W}{w}\times\frac{H}{h}$ in the real world. The BEV query is a 3D trainable embedding with a dimension of $l\times w\times h$ representing the BEV space, and serving as the query for deformable attention modules in the encoder. All intermediate BEV features in the network also follow the same spatial dimension. The specific range and dimension chosen for our experiment are described in Sec. IV-A.
@@
$\mathcal{X}^{*}=\operatorname*{arg\,min}_{\mathcal{X}}\psi\Big{(}I^{{}^{\prime}}_{\text{bev}}(\mathcal{X}),I_{\text{map}}(\mathcal{X})\Big{)},$ (2)
where $\psi$ is a function to find $\mathcal{X}^{*}$ to achieve minimum distance between two representations, and provided by template matching in our system. $I_{\text{map}}$ is the subset of aerial map with respect to vehicle pose, and $I^{{}^{\prime}}_{\text{feat}}$ is the rendered BEV image given learned feature:
$I^{{}^{\prime}}_{\text{bev}}(\mathcal{X})=\varphi_{\text{render}}\Big{(}F_{\text{feat}}(\mathcal{X})\Big{)},$ (3)
where $\varphi_{\text{render}}$ is a mapping from the encoded feature $F_{\text{feat}}$ to top-down BEV image, given by the rendering head.
@@ III-A Feature Encoding with BEVFormer
Adopting BEVFormer’s framework [2], we propagate consecutive frame features to capture temporal information. Within a temporal window of $T$ seconds, $n$ frames ($3\times n$ images in a trinocular setup) are sampled. A detailed setting can be found in Sec. IV-A. Starting with the earliest frames, camera images $I^{\mathrm{cam}}_{t}$ are processed through patch projection, which is a convolutional layer in our implementation, to obtain camera feature $F^{\mathrm{cam}}_{t}$ and sent to the encoder together with the BEV query $Q$ and previous BEV feature $B_{t-1}$ to obtain the encoded BEV feature for current timestamp $B_{t}$. The encoding process consists of two stages: a temporal attention stage that takes in query $Q$ and previous timestamp BEV feature $B_{t-1}$ for deformable attention:
@@
$B^{\text{spatial}}_{t}=\text{DeformableAttn}\Big{(}F^{\text{cam}}_{t},B^{\text{temp}}_{t}\Big{)},$ (5)
$B_{t}$ is then projected to the location of the subsequent frame as $B^{{}^{\prime}}_{t}$ according to the movement of the vehicle given by the GPS information, using affine transformations in $\mathbf{SE}(2)$ and bilinear interpolation:
$\Delta\mathcal{X}=\Big{[}\Delta x,\Delta y,\Delta\theta\Big{]}=\mathcal{X}_{t}-\mathcal{X}_{t-1},$ (6)
@@
Subsequently, $B^{{}^{\prime}}_{t}$ serves as a query to the encoder together with the next camera feature $F_{t+1}$ to obtain $B_{t+1}$. Propagation continues in the temporal window until we obtain the latest timestamp feature $B^{T}$. A diagram of temporal propagation is shown in Fig. 3. It should be noted that $B_{t-1}$ is the same as query $Q$ for the first frame in the temporal window:
@@ IV-A Experiment Setting
During the testing phase, we crop and rotate the aerial map based on the GPS ground truth position as the center of the image with a size of 874$\times$874 pixels, which corresponds to a real-world coverage of approximately 200$\times$200 square meters. This search region is sufficient to accommodate VIO drift for more than 10 minutes without a registration. For cross-sequence testing, loosen the assumption of drifting range and use a search region of 100$\times$100 square meters. Our camera system captures 3 frames per second and predicts registration consistent with camera frame; therefore, sufficient for preventing from failing with the 100$\times$100 square meter search range.
We use NCC for template matching. NCC identifies the best match within the search area, maximizing similarity between the generated BEV image and the aerial map, thus predicting the vehicle’s position relative to the aerial map. We observe failure cases when rendered BEV images are of moderate visual quality, whereas NCC fails in prediction. An example of failure cases is shown in Fig. 6.
@@ IV-B Dataset Organization
We collect our real-world data set in the Pittsburgh area, with a VIO system on board. Detailed information on the sequences can be found in Table II. For each training sample, we use the information of timestamp, trinocular RGB images, and GPS ground truth including x, y, and azimuth angle in the UTM coordinate system for training. The pre-processing process for cropping the aerial map can be found in Fig. 1.
@@ IV-C Quantitative Comparison
$\mathbf{p}_{i}$ $=(x_{i},y_{i}),$ (11)
$\mathbf{d}^{i}_{\mathrm{Euclidean}}$ $=\lVert\mathbf{p}^{i}_{\mathrm{gt}}-\mathbf{p}^{i}_{\mathrm{{pred}}}\rVert_{2},$ (12)
$p_{\mathrm{match}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\bm{\cdot}(\mathbf{d}_{\mathrm{Euclidean}}^{i}<\mathbf{d}_{\mathrm{threshold}}),$ (13)
@@ Acknowledgement
The authors thank Kaicheng Yu and Wenshan Wang for the fruitful discussions and valuable suggestions throughout the project, and staff members of the National Robotics Engineering Center (NREC) for helping with data collection.

### v2

@@ Abstract
We introduce BEVRender, a novel learning-based approach for the localization of ground vehicles in Global Navigation Satellite System (GNSS)-denied off-road scenarios. These environments are typically challenging for conventional vision-based state estimation due to the lack of distinct visual landmarks and the instability of vehicle poses. To address this, BEVRender generates high-quality local bird’s-eye-view (BEV) images of the local terrain. Subsequently, these images are aligned with a georeferenced aerial map through template matching to achieve accurate cross-view registration. Our approach overcomes the inherent limitations of visual inertial odometry systems and the substantial storage requirements of image-retrieval localization strategies, which are susceptible to drift and scalability issues, respectively. Extensive experimentation validates BEVRender’s advancement over existing GNSS-denied visual localization methods, demonstrating notable enhancements in both localization accuracy and update frequency.
@@ III Methodology
Our system contains three main components: a feature encoder to map the visual features from the camera to the top-down view, a rendering head to decode the features and generate top-down BEV images, and an image registration component for cross-view localization. An overview of our system is shown in Fig. 1.
@@
$\mathbf{X}_{t}=\Big{[}x_{t},y_{t},\theta_{t}\Big{]}\in\mathbf{SE}(2).$ (1)
We assume that the prediction for the azimuth angle $\theta_{t}$ from VIO is accurate, but the position estimates ($x_{t}$ and $y_{t}$) may drift over time. Our system aggregates information from consecutive frames to construct a top-down representation of the environment for map registration.
We define a 3D BEV space centered on the vehicle with a length of $L$, a width of $W$ and a height of $H$. The space is divided into $l\times w\times h$ grid cells, so that each cell represents a cubic size of $\frac{L}{l}\times\frac{W}{w}\times\frac{H}{h}$ in the real world. The BEV query is a 3D trainable embedding with a dimension of $l\times w\times h$ representing the BEV space and serving as the query for deformable attention modules in the encoder. All intermediate BEV features in the network also follow the same spatial dimension. The specific range and dimension chosen for our experiment are described in Sec. IV-A.
@@
$\mathbf{X}^{*}=\operatorname*{arg\,min}_{\mathbf{X}}\Phi\Big{(}I^{{}^{\prime}}_{\text{bev}}(\mathbf{X}),I_{\text{map}}(\mathbf{X})\Big{)},$ (2)
where $\Phi$ is a function to find $\mathbf{X}^{*}$ to achieve minimum distance between two representations in feature space, and is provided by template matching module in our system. $I_{\text{map}}$ is a subset of the aerial map with respect to a vehicle pose.
The image rendering head $\Psi_{\text{render}}$ defines a mapping from the encoded image feature $F_{\text{feat}}$ to top-down BEV image $I^{{}^{\prime}}_{\text{bev}}$:
$I^{{}^{\prime}}_{\text{bev}}(\mathbf{X})=\Psi_{\text{render}}\Big{(}F_{\text{feat}}(\mathbf{X})\Big{)}.$ (3)
@@ III-A Feature Encoding with BEVFormer
Adopting BEVFormer’s framework [2], we propagate consecutive frame features to capture temporal information. Within a temporal window of $T$ seconds, $n$ frames ($3\times n$ images in a trinocular setup) are sampled. A detailed setting can be found in Sec. IV-A. Starting with the earliest frame, camera images $I^{\mathrm{cam}}_{t}$ are processed through patch projection, which is a convolutional layer in our implementation, to obtain camera feature $F^{\mathrm{cam}}_{t}$ and sent to the encoder together with the BEV query $Q$ and previous BEV feature $B_{t-1}$ to obtain the encoded BEV feature for current timestamp $B_{t}$. The encoding process consists of two stages: a temporal attention stage that takes in query $Q$ and previous timestamp BEV feature $B_{t-1}$ for deformable attention:
@@
$B_{t}=B^{\text{spatial}}_{t}=\text{DeformableAttn}\Big{(}F^{\text{cam}}_{t},B^{\text{temp}}_{t}\Big{)}.$ (5)
$B_{t}$ is then projected to the subsequent frame vehicle pose as $B^{{}^{\prime}}_{t}$ according to the movement of the vehicle provided by GPS. The projection is performed by affine transformations in $\mathbf{SE}(2)$ followed by a bilinear interpolation:
$\Delta\mathbf{X}=\mathbf{X}_{t}-\mathbf{X}_{t-1}=\Big{[}\Delta x,\Delta y,\Delta\theta\Big{]},$ (6)
@@
Subsequently, $B^{{}^{\prime}}_{t}$ serves as a query to the encoder to query the next timestamp camera feature $F_{t+1}$ to obtain $B_{t+1}$. The propagation continues in the temporal window until we obtain the latest timestamp feature. A diagram of temporal propagation is shown in Fig. 3. It should be noted that $B_{t-1}$ is the same as query $Q$ for the first frame in the temporal window:
@@ IV-A Experiment Setting
During the testing phase, we crop and rotate the aerial map based on the GPS ground truth position as the center of the image with a size of 874$\times$874 pixels, which corresponds to a real-world coverage of approximately 200$\times$200 square meters. This search region is sufficient to accommodate VIO drift for more than 10 minutes without registration. For cross-sequence testing, we loosen the assumption of drifting range and use a search region of 100$\times$100 square meters. Our camera system captures 3 frames per second and predicts registration consistent with camera frame; therefore, sufficient for preventing from failing with the 100$\times$100 square meter search range.
We use NCC for template matching. NCC identifies the best match within the search area, maximizing similarity between the generated BEV image and the aerial map, thus predicting the vehicle’s position relative to the aerial map. We observe failure cases where rendered BEV images are of moderate visual quality, whereas NCC fails in prediction. An example of failure cases is shown in Fig. 6.
@@ IV-B Dataset Organization
We collect our real-world data set in the Pittsburgh area, with a VIO system on board. Detailed information on the sequences can be found in Table II. For each training sample, we use the information of timestamp, trinocular RGB images, and GPS ground truth including x, y, and azimuth angle in the UTM coordinate system for training. The preprocessing for cropping the aerial map can be found in Fig. 1.
@@ IV-C Quantitative Comparison
$\mathbf{x}^{i}$ $=(x_{i},y_{i}),$ (11)
$\mathbf{d}^{i}$ $=\lVert\mathbf{x}^{i}_{\mathrm{gt}}-\mathbf{x}^{i}_{\mathrm{{pred}}}\rVert_{2},$ (12)
$\mathbf{p}_{\mathrm{match}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\bm{\cdot}(\mathbf{d}^{i}<\mathbf{d}_{\mathrm{threshold}}),$ (13)
@@ Acknowledgement
The authors thank Kaicheng Yu for the fruitful discussion and valuable suggestions, and members of the National Robotics Engineering Center (NREC) for helping with data collection.